OpenStudy (anonymous):

CALC QUESTION: (Will medal) An object moves along the x-axis with velocity of v(t) = t(sqrt(2t+3)) -3t For which 't' is the object moving to the right?

OpenStudy (anonymous):

Will I have to integrate?

OpenStudy (anonymous):

\[v(t) = t \sqrt{2t+3} - 3t\]

OpenStudy (anonymous):

\[\int\limits_{?}^{?}t \sqrt{2t+3} - 3t?????\]

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi!

OpenStudy (misty1212):

no \(v\) is the velocity, same as speed except it can be positive or negative you are being asked where (for what value of \(t\) ) is it positive

OpenStudy (anonymous):

Oh so I use the acceleration I found (second derivative)? Where would I go from there?

OpenStudy (misty1212):

you are thinking too hard

OpenStudy (misty1212):

if the velocity is positive, it is going to the right

OpenStudy (anonymous):

Should I plug in 0 to the velocity function?

OpenStudy (misty1212):

\[t\sqrt{2t+3}-3t>0\] solve for \(t\)

OpenStudy (anonymous):

Oh ok!

OpenStudy (anonymous):

I got t > 3. Thank you! (Hope this is right) ((I need to be more confident).